# Bilevel Learning for Bilevel Planning

### v2

@@ IV-A Effect Vectors as Supervision for Neural Learning
Figure 3: Detailed algorithm for inventing predicate P2(?r,?t). At the $t$-th iteration, the symbolic-level proposes $\Delta^{\psi}_{t}=[+1,+1,0,0]$ (right bottom), which is used to derive supervisions for the two ground transition pairs. Due to the unreasonable effects, the intermediate state is labeled as both True and False, resulting in high validation loss $\mathbf{L}_{t}$. The loss then informs the symbolic learning in the $t+1$-th iteration.
@@
For example, in Figure 3, the effect vector $[+1,+1,0,0]$ specifies that predicate $\psi=\mathtt{P2\_1(?r,?t)}$ is the add effect for both $\mathtt{Gaze(?r,?t)}$ and $\mathtt{Grasp(?r,?t)}$ [^2] Each predicate here can appear at most once in the effect sets, but this doesn’t affect the representation capability of the final predicate set..
Lifted effect vector is a favorable symbolic representation of a lifted predicate, since its shape doesn’t depend on task object compositions and can thus be learned more efficiently, as we will see. However, to train the neural classifier on the transition pairs, we will need to derive some supervisions on ground predicates, which is achieved by ground effect vector.
@@
where $\underline{\psi}_{p}$ denotes the $p$-th atom with objects $\mathcal{O}_{\underline{\psi}_{p}}$, among the total of $P$ ground predicates. For example, in Figure 3, ground effects for $\mathtt{P2\_1(r_{1},t_{1})}$ will be $+1$ for both ground actions $\mathtt{Gaze(r_{1},t_{1})}$ and $\mathtt{Grasp(r_{1},t_{1})}$, while ground effects for $\mathtt{P2\_1(r_{1},t_{2})}$ will be 00, as $\mathtt{(r_{1},t_{2})}\not\subseteq\mathtt{(r_{1},t_{1})}$.
Importantly, this implies that the “value” of the (potentially) non-zero entry of all ground effects from the action class $\mathtt{C}_{i}$ equals $\delta^{\psi}_{i}$, while the “position” of the non-zero entry is decided by the object set $\mathcal{O}_{\underline{\mathtt{C}}}$ and the predicate grounding (more explanations and discussions see Appendix -B and Appendix -C).
@@
where $\hat{\mathbf{v}},\hat{\mathbf{v}}^{\prime}\in[0,1]^{P}$ are the predicted ground predicates by applying $\theta_{\psi}$ on all possible object sets from $\mathcal{O}$. $\mathrm{Div}_{\mathrm{JS}}(\cdot|\cdot)$ denotes the Jensen–Shannon divergence [34] and Eq.(3) tries to minimize the distance between $\hat{\mathbf{v}}$ and $\hat{\mathbf{v}}^{\prime}$ if the indices with zero values in $\bm{t}^{\psi,\underline{\mathtt{C}}}$. $\hat{{v}}_{p},\hat{{v}}_{p}^{\prime}$ denotes $p$-th ground predicate, where $\mathcal{O}_{\underline{\psi}_{p}}\subseteq\mathcal{O}_{\underline{\mathtt{C}}_{i}}$. $\mathrm{BCE}(\cdot,\cdot)$ represents the Binary Cross-Entropy, which tries to directly supervise $\hat{\mathbf{v}}$ and $\hat{\mathbf{v}}^{\prime}$ if $\delta^{\psi}_{i}\neq 0$. Since we have $\Delta^{\psi}$ for all lifted actions, the pipeline can be applied to all ground transition pairs in $\mathcal{D}$, resulting in the global loss,
@@
where $\mathcal{D}_{\mathtt{C}}$ denotes the distribution of the grounded transition for action $\mathtt{C}$ in the dataset $\mathcal{D}$ (See Figure 3 for the examples of two transitions from different actions). Since $\mathcal{L}$ is fully differentiable with respect to $\theta$, given a effector $\Delta^{\psi}$ and the demonstration dataset, we could leverage the general and standard deep learning framework [35, 36] to train a neural classifier $\theta$ that minimizes the loss $\mathcal{L}$ for any state representation.
@@ IV-B Neural Loss as Guidance for Symbolic Learning
To obtain all the classifiers $\theta_{\Psi^{\mathtt{Var}}}$ for all the typed predicates $\psi\in\Psi^{\mathtt{Var}}$, the problem now becomes finding all the lifted effect vectors $\Delta^{\psi}\in\Delta^{\Psi^{\mathtt{Var}}}$. As defined in Definition 3, the lifted effect vectors live in the discrete world with a finite shape, which motivates us to establish some discrete optimization strategies for it. One insight from IVNTR is that, despite the large space, only a few effect vectors provide reasonable supervision, with unreasonable ones resulting in high classification error on the validation set after the training converges.
@@
IV-C Predicate Selection
With all possible typed predicates, IVNTR is able to construct the predicate pool $\hat{\Psi}_{\mathrm{dyn}}$ without any function pre-definition. This strength has made bilevel planning applicable to complicated and high-dimensional state spaces. Meanwhile, as the predicate functions are relational and can be seamlessly integrated into an AI planner [33], our approach can naturally achieve compositional generalization.
Yet, not all of the predicates in $\hat{\Psi}_{\mathrm{dyn}}$ are favorable for planning. Therefore, we next try to select a subset $\tilde{\Psi}_{\mathrm{dyn}}\subset\hat{\Psi}_{\mathrm{dyn}}$ that minimizes the score function $J(\tilde{\Psi}_{\mathrm{dyn}},\mathcal{D})$ [16]. Specifically, with a set of candidate predicates $\tilde{\Psi}=\{\Psi_{\mathrm{G}},\Psi_{\mathrm{sta}},\tilde{\Psi}_{\mathrm{dyn}}\}$, we start by grounding all the states $\mathbf{x}$ in the demonstration $\mathcal{D}$, which forms a ground atom dataset. Since we already have the effect vector for each of the predicates in $\tilde{\Psi}_{\mathrm{dyn}}$, the effect sets ($\tilde{\mathtt{Eff}}^{+},\tilde{\mathtt{Eff}}^{-}\subset\tilde{\Psi}_{\mathrm{dyn}}$) for each operator $\tilde{\mathtt{Op}}^{\mathtt{C}}$ can be easily obtained. Then, the precondition set for each operator can be calculated using an intersection strategy [31]. The learned operators set $\tilde{\mathtt{Op}}$ is then applied to the ground atom dataset to generate plan skeletons $\tilde{\pi}$ for each task, which are compared with the demonstration plan skeletons $\bar{\pi}$ for objective calculation [16]. The objective is finally minimized by running a hill-climbing search over $\hat{\Psi}_{\mathrm{dyn}}$, resulting in the desired predicate set ${\Psi}_{\mathrm{dyn}}$. With the complete set, we are now able to learn the planning abstractions (operators and samplers) using standard pipelines [31, 19, 20] as shown in Figure 2 (a).
Note that before the predicate selection stage, all of the predicates’ neural functions have been pre-trained using our IVNTR framework, which has avoided the challenge of using the planning objective for learning but still achieved a powerful and adaptive model optimized for planning.
@@ V-A Implementation Details
TABLE III: Comparison between the predicates learning using ground-truth (GT)-vectors and using our IVNTR framework. We report the task success rate and the final planning objective.
@@
V-B Empirical Results
Simulated Planning Domains:  The empirical comparison across the six simulated domains is presented in Table I. Alongside the averaged success rates, we report the performance drop percentage during generalization. IVNTR consistently outperforms all baselines in both $\mathcal{T}^{\mathrm{train}}$ and $\mathcal{T}^{\mathrm{test}}$ across all domains. For complex state representations such as SE3 and PointClouds (PCD), none of the baselines achieve a success rate above $5\%$ on generalized tasks, while IVNTR stably solves over $50\%$ by virtue of its relational structure.
V-C Ablation Studies
Comparison to Other Search Algorithms:  To evaluate the efficiency of our neural-informed tree expansion algorithm, we compared it with alternative search strategies: a greedy approach (Greedy) that flips the zero entry with the highest current loss, breadth-first (BFS) and depth-first (DFS) searches. For the Engrave domain, Figure 5 shows the number of iterations required to find each reasonable effect vector using these methods. Compared to uninformed methods, our IVNTR framework is at least 2×2$\times$2 × more efficient. The greedy approach exhibits high variance and is generally less reliable.
V-D Interpreting Invented Predicates
@@
TABLE IV: Comparison between sampling with and without the invented predicate functions. Without using the functions as step-wise success indicator, the performance drops significantly.
Predicates play a key role in defining the preconditions and effects of operators, specifying the order of ground actions to complete a task. In Table V, we display part of the precondition and effect sets for high-level actions in the Climb-Transport domain, where invented predicates capture logical relationships among actions. For example, the $\mathtt{Drop}$ action requires $\mathtt{P2\_1}$ and $\mathtt{P2\_2}$, which are the add effects of $\mathtt{Gaze}$, $\mathtt{MTGaze}$, and $\mathtt{WalkOn}$. Similarly, the $\mathtt{Pick}$ action depends on $\mathtt{P1\_1}$, the delete effect of $\mathtt{Grasp}$, enforcing all $\mathtt{Pick}$ actions to precede $\mathtt{Grasp}$. These relational constraints over objects enable the generation of long-horizon plans that generalize to unseen object compositions. The complete operators are detailed in Appendix -H. We also provide visualizations of how predicates act as success indicators to filter out sampler failures in Appendix -I.
@@ VI-B Task and Motion Planning
Predicates $\mathtt{P1\_1(?r)}$ $\mathtt{P2\_1(?r,?t)}$ $\mathtt{P2\_2(?r,?t)}$ $\mathtt{In(?t,?t)}$
@@ VI-C Predicate Invention for Planning
To automate predicate generation for planning, various approaches have been proposed [20, 23, 24, 16, 27, 28, 26, 29]. The most direct approaches [23, 24] rely on domain knowledge, such as human-provided labels [23] or LLM-based oracles [24]. To invent predicates, earlier approaches derive ”easy” step-wise objectives, such as reconstruction [26, 27, 28] or bisimulation [29]. Among these, LatPlan (FOSAE) [27] learns relational neural abstractions for images by reconstructing states and identifying action spaces for planning, which is the closest work to IVNTR. However, its implicit predicates are not optimized for efficient planning, limiting its applicability to domains with only nullary actions. Recent approaches[16, 20] address this by learning abstractions tailored to fast planners [33]. However, these methods struggle to learn predicate functions due to non-differentiable objectives. Consequently, they often rely on pre-defined predicate candidates from program synthesis [16] or foundation models [20], which constrains their applicability in more sophisticated and high-dimensional state spaces. Our approach is motivated by the bilevel planning framework [16] but eliminates the need for pre-defining the candidates. Instead, we can learn adaptive neural functions for different domains, enabling more flexible and scalable learning based planning.
@@ VII Limitations and Future Works
In this work, we introduced IVNTR, a bilevel learning framework that invents neural classifiers as relational planning predicates. These predicates enable the learning of relational bilevel planners capable of generating long-horizon plans for unseen object compositions. At the neural level, IVNTR leverages the high-level effects of predicates across actions to provide step-wise supervisions on transition pairs. At the symbolic level, IVNTR captures the sparsity of effects through an informed tree expansion algorithm. By adopting neural functions, IVNTR adapts to diverse robot planning domains with continuous and high-dimensional state representations. Additionally, we deployed IVNTR on a mobile manipulator, demonstrating its ability to achieve compositional generalization over objects and actions in long-horizon mobile manipulation tasks.
IVNTR has several limitations that we leave for future work: (1) IVNTR can only invent dynamic predicates. Static predicates are still assumed as domain-level prior knowledge. (2) Following previous works [18], we have assumed the sparsity of effects; discussions about the general cases can be found in Appendix -B. (3) Since IVNTR trains different neural networks in each iteration, the learning time could be prolonged in domains with extreme complexity. Currently, we parallelize the invention of different predicate groups on multiple GPUs for efficiency (see Appendix -E). (4) Neural predicates with quantifiers are less reliability due to the errors in neural classifiers. Future work could explore the probabilistic symbolic planners [61] to handle the neural grounding noise.

### v3

@@ IV-A Effect Vectors as Supervision for Neural Learning
Figure 3: Detailed neural learning process for predicate $\mathtt{P2_{1}(?r,?t)}$. From the demonstration dataset, we display two transition pairs (one for each action, in total four states) on the left. The neural network takes object centric features as input, predicting ground predicates (in total eight values). At the bottom, we display an example lifted effect vector for action Grasp, Gaze as $\Delta^{\psi}_{t}=[+1,+1]$. With the ground effect vector, supervisions can be derived on the predicted values. Due to the unreasonable effect supervisions, the intermediate state is labeled as both True and False, resulting in high validation loss.
@@
For example, in Figure 3, the effect vector $[+1,+1]$ specifies that predicate $\psi=\mathtt{P2_{1}(?r,?t)}$ is the add effect for both $\mathtt{Gaze(?r,?t)}$ and $\mathtt{Grasp(?r,?t)}$ [^2] Each predicate here can appear at most once in the effect sets, but this doesn’t affect the representation capability of the final predicate set..
The lifted effect vector is a favorable symbolic representation of a lifted predicate, since its shape doesn’t depend on task object compositions and can thus be learned more efficiently, as we will see. However, to train the neural classifier on the transition pairs, we will need to derive supervision on ground predicates, which is achieved by the ground effect vector.
@@
where $\underline{\psi}_{p}$ denotes the $p$-th atom with objects $\mathcal{O}_{\underline{\psi}_{p}}$, among the total of $P$ ground predicates. For example, in Figure 3, ground effects for $\mathtt{P2_{1}(r_{1},t_{1})}$ will be $+1$ for both ground actions $\mathtt{Gaze(r_{1},t_{1})}$ and $\mathtt{Grasp(r_{1},t_{1})}$, while ground effects for $\mathtt{P2_{1}(r_{1},t_{2})}$ will be 00, as $\mathtt{(r_{1},t_{2})}\not\subseteq\mathtt{(r_{1},t_{1})}$.
Importantly, this implies that the “value” of the (potentially) non-zero entry of all ground effects from the action class $\mathtt{C}_{i}$ equals $\delta^{\psi}_{i}$, while the “position” of the non-zero entry is decided by the object set $\mathcal{O}_{\underline{\mathtt{C}}}$ and the predicate grounding (see Appendix -B and Appendix -C for more explanations).
@@
where $\hat{\mathbf{v}},\hat{\mathbf{v}}^{\prime}\in[0,1]^{P}$ are the predicted ground predicates by applying $\theta_{\psi}$ on all possible object sets from $\mathcal{O}$. $\mathrm{Div}_{\mathrm{JS}}(\cdot|\cdot)$ denotes the Jensen–Shannon divergence [35] and Eq.(3) tries to minimize the distance between $\hat{\mathbf{v}}$ and $\hat{\mathbf{v}}^{\prime}$ if the indices with zero values in $\bm{t}^{\psi,\underline{\mathtt{C}}}$. $\hat{{v}}_{p},\hat{{v}}_{p}^{\prime}$ denotes $p$-th ground predicate, where $\mathcal{O}_{\underline{\psi}_{p}}\subseteq\mathcal{O}_{\underline{\mathtt{C}}_{i}}$. $\mathrm{BCE}(\cdot,\cdot)$ represents the Binary Cross-Entropy, which tries to directly supervise $\hat{\mathbf{v}}$ and $\hat{\mathbf{v}}^{\prime}$ if $\delta^{\psi}_{i}\neq 0$. Intuitively, $\mathcal{L}_{\mathrm{zero}}$ supervises the ground predicates whose values should stay unchanged in a transition (but we don’t know their values). $\mathcal{L}_{\mathrm{one}}$, on the other hand, supervises the ground predicates whose values can be derived based on the effect vectors. Since we have $\Delta^{\psi}$ for all lifted actions, the pipeline can be applied to all ground transition pairs in $\mathcal{D}$, resulting in the global loss,
@@
where $\mathcal{D}_{\mathtt{C}}$ denotes the distribution of the grounded transition for action $\mathtt{C}$ in the dataset $\mathcal{D}$ (See Figure 3 for the examples of two transitions from different actions). Since $\mathcal{L}$ is fully differentiable with respect to $\theta$, given a effector $\Delta^{\psi}$ and the demonstration dataset, we could leverage the general and standard deep learning frameworks [36, 37] to train a neural classifier $\theta$ that minimizes the loss $\mathcal{L}$ for any state representation.
@@ IV-B Neural Loss as Guidance for Symbolic Learning
To obtain all the classifiers $\theta_{\Psi^{\mathtt{Var}}}$ for all the typed predicates $\psi\in\Psi^{\mathtt{Var}}$, the problem now becomes finding all the lifted effect vectors $\Delta^{\psi}\in\Delta^{\Psi^{\mathtt{Var}}}$. As defined in Definition 3, the lifted effect vectors live in the discrete world with a finite shape, which motivates us to establish some discrete optimization strategies for it. One insight here is that, despite the large space, only a few effect vectors provide reasonable supervision, with unreasonable ones resulting in high classification error on the validation set after the training converges.
@@
Figure 4: Detailed symbolic learning process for predicate group P2(?r,?t). With the neural validation loss from the previous iteration, symbolic learning starts by merging the loss into the global value vector (See Eq. (7)). After the node values in the effect tree are updated, we conduct parent node selection and expansion. Among the children nodes, we evaluate the child with the current highest value, via the neural classifer training process described in Figure 3.
@@
IV-C Predicate Selection
With all possible typed predicates, IVNTR is able to construct the predicate pool $\hat{\Psi}_{\mathrm{dyn}}$ without any classifier pre-definition. This strength has made bilevel planning applicable to complicated and high-dimensional state spaces. Meanwhile, as the predicate classifiers are relational and can be seamlessly integrated into an AI planner [34], our approach can naturally achieve compositional generalization.
Yet, not all of the predicates in $\hat{\Psi}_{\mathrm{dyn}}$ are favorable for planning. Therefore, we next try to select a subset $\tilde{\Psi}_{\mathrm{dyn}}\subset\hat{\Psi}_{\mathrm{dyn}}$ that minimizes the score function $J(\tilde{\Psi}_{\mathrm{dyn}},\mathcal{D})$ [17]. Specifically, with a set of candidate predicates $\tilde{\Psi}=\{\Psi_{\mathrm{G}},\Psi_{\mathrm{sta}},\tilde{\Psi}_{\mathrm{dyn}}\}$, we start by grounding all the states $\mathbf{x}$ in the demonstration $\mathcal{D}$, which forms a ground atom dataset. Since we already have the effect vector for each of the predicates in $\tilde{\Psi}_{\mathrm{dyn}}$, the effect sets ($\tilde{\mathtt{Eff}}^{+},\tilde{\mathtt{Eff}}^{-}\subset\tilde{\Psi}_{\mathrm{dyn}}$) for each operator $\tilde{\mathtt{Op}}^{\mathtt{C}}$ can be easily obtained. Then, the precondition set for each operator can be calculated using an intersection strategy [32]. The learned operator set $\tilde{\mathtt{Op}}$ is then applied to the ground atom dataset to generate plan skeletons $\tilde{\pi}$ for each task, which are compared with the demonstration plan skeletons $\bar{\pi}$ for objective calculation [17]. The objective is finally minimized by running a hill-climbing search over $\hat{\Psi}_{\mathrm{dyn}}$, resulting in the desired predicate set ${\Psi}_{\mathrm{dyn}}$. With the complete set, we are now able to learn the planning abstractions (operators and samplers) using standard pipelines [32, 20, 21] as shown in Figure 2 (a). For a more detailed explanation to operator and sampler learning, please see Appendix Appendix -K.
Note that before the predicate selection stage, all of the predicates’ neural classifiers have been pre-trained using our IVNTR framework, which has avoided the challenge of using the planning objective for learning but still achieved a powerful and adaptive model optimized for planning.
@@ V Experiments
TABLE III: Comparison between the predicates learned with ground-truth (GT)-vectors and using our IVNTR framework. We report the task success rate and the final planning objective.
V-A Implementation Details
@@
TABLE IV: Comparison between sampling with and without the invented predicate classifiers. Without using the predicates as step-wise success indicator, the performance drops significantly.
V-B Empirical Results
Simulated Planning Domains:  The empirical comparison across the six simulated domains is presented in Table I. Alongside the averaged success rates, we report the performance drop percentage during generalization. IVNTR consistently outperforms all baselines in both $\mathcal{T}^{\mathrm{train}}$ and $\mathcal{T}^{\mathrm{test}}$ across all domains. For complex state representations such as SE3 and PointClouds (PCD), none of the baselines achieve a success rate above $5\%$ on generalized tasks, while IVNTR stably solves over $50\%$ by virtue of its relational structure. In Appendix -L, we further demonstrate the potential of IVNTR to abstract high-dimensional RGB image states into symbolic predicates.
V-C Ablation Studies
Comparison to Other Search Algorithms:  To evaluate the efficiency of our neural-informed tree expansion algorithm, we compared it with alternative search strategies: a greedy approach (Greedy) that flips the zero entry with the highest current loss, breadth-first (BFS) and depth-first (DFS) searches. For the Engrave domain, Figure 6 shows the number of iterations required to find each reasonable effect vector using these methods. Compared to uninformed methods, our IVNTR framework is at least 2×2$\times$2 × more efficient. The greedy approach exhibits high variance and is generally less reliable. We present comparisons in more domains in Appendix -M.
V-D Interpreting Invented Predicates
Predicates play a key role in defining the preconditions and effects of operators, specifying the order of ground actions to complete a task. In Table V, we display part of the precondition and effect sets for high-level actions in the Climb-Transport domain, where invented predicates capture logical relationships among actions. For example, the $\mathtt{Drop}$ action requires $\mathtt{P2_{1}}$ and $\mathtt{P2_{2}}$, which are the add effects of $\mathtt{Gaze}$, $\mathtt{MTGaze}$, and $\mathtt{WalkOn}$. Similarly, the $\mathtt{Pick}$ action depends on $\mathtt{P1_{1}}$, the delete effect of $\mathtt{Grasp}$, enforcing all $\mathtt{Pick}$ actions to precede $\mathtt{Grasp}$. These relational constraints over objects enable the generation of long-horizon plans that generalize to unseen object compositions. The complete operators are detailed in Appendix -H. We also provide visualizations of how predicates act as success indicators to filter out sampler failures in Appendix -I.
@@ VI-B Task and Motion Planning
Predicates $\mathtt{P1_{1}(?r)}$ $\mathtt{P2_{1}(?r,?t)}$ $\mathtt{P2_{2}(?r,?t)}$ $\mathtt{In(?t,?t)}$
@@ VI-C Predicate Invention for Planning
To automate predicate generation for planning, various approaches have been proposed [21, 24, 25, 17, 28, 29, 27, 30, 14]. The most direct approaches [24, 25] rely on domain knowledge, such as human-provided labels [24] or LLM-based oracles [25]. To invent predicates, earlier approaches derive ”easy” step-wise objectives, such as reconstruction [27, 28, 29] or bisimulation [30]. Among these, LatPlan (FOSAE) [28] learns relational neural abstractions for images by reconstructing states and identifying action spaces for planning, which is the closest work to IVNTR. However, its implicit predicates are not optimized for efficient planning, limiting its applicability to domains with only nullary actions. Recent approaches [17, 21] address this by learning abstractions tailored to fast planners [34]. However, these methods struggle to learn predicate classifiers due to non-differentiable objectives. Consequently, they often rely on pre-defined predicate candidates from program synthesis [17] or foundation models [21], which constrains their applicability in more sophisticated and high-dimensional state spaces. Our approach is motivated by the bilevel planning framework [17] but eliminates the need for pre-defining the candidates. Instead, we can learn adaptive neural classifiers for different domains, enabling more flexible and scalable learning based planning.
@@ VII Limitations and Future Works
In this work, we introduced IVNTR, a bilevel learning framework that invents neural classifiers as relational planning predicates. These predicates enable the learning of relational bilevel planners capable of generating long-horizon plans for unseen object compositions. At the neural level, IVNTR leverages the high-level effects of predicates across actions to provide step-wise supervisions on transition pairs. At the symbolic level, IVNTR captures the sparsity of effects through an informed tree expansion algorithm. By adopting neural classifiers, IVNTR adapts to diverse robot planning domains with continuous and high-dimensional state representations. Additionally, we deployed IVNTR on a mobile manipulator, demonstrating its ability to achieve compositional generalization over objects and actions in long-horizon mobile manipulation tasks.
IVNTR has several limitations that we leave for future work: (1) IVNTR can only invent dynamic predicates. Static predicates are still assumed as domain-level prior knowledge. (2) Following previous works [19], we have assumed the sparsity of effects; discussions about the general cases can be found in Appendix -B. (3) Since IVNTR trains different neural networks in each iteration, the learning time could be prolonged in domains with extreme complexity. Currently, we parallelize the invention of different predicate groups on multiple GPUs for efficiency (see Appendix -E). (4) Neural predicates with quantifiers are less reliable due to the prediction errors in the classifiers. Future work could explore probabilistic symbolic planners [62] or hybrid declarative-imperative representations [63] to plan with an inaccurate model. (5) Since the predicates are neural networks, it is hard to interpret their physical meanings. It would also be intriguing for future approaches to make them directly human readable.